# The role of spatial dimension in the emergence of localised radial patterns from a Turing instability

### v1

@@ 3.2 The Far-Field Manifold
We now turn to parametrising the far-field manifold $\mathcal{W}_{+}^{s}(\mu)$, consisting of all small-amplitude solutions to (2.1) that decay to zero exponentially fast as $r\to\infty$. We introduce the variable $\sigma(r)\geq 0$ to take the place of the $\frac{1}{r}$ terms in (3.1). The result is the extended autonomous system
@@
with the property that $\sigma(r)=r^{-1}$ is an invariant manifold of (3.18), and by definition this invariant manifold recovers the non-autonomous system (3.1). To construct the far-field manifold $\mathcal{W}^{s}_{+}(\mu)$ we find the set of all small-amplitude solutions $(\mathbf{U},\sigma)(r)$ to (3.18) such that $\mathbf{U}(r)$ decays exponentially as $r\to\infty$. Then, evaluating at $\sigma(r_{0})=r_{0}^{-1}$, we restrict our solutions to the invariant subspace $\sigma(r)=r^{-1}$, such that $\mathbf{U}(r)$ is an exponentially decaying solution of (3.1).
Before attempting to find exponentially decaying solutions to (3.18), we first transform the system into the normal form for an $n$-dimensional Hamiltonian–Hopf bifurcation. We define complex amplitudes $\widetilde{A},\widetilde{B}$, which satisfy the following relations
@@
In the coordinates $\widetilde{A},\widetilde{B}$, defined by (3.19), the core manifold (3.17) can be expressed as
@@
where we recall that $\nu_{n}=\left(\frac{3}{8}\right)^{\frac{n}{2}}\frac{\pi}{3\Gamma(\frac{n}{2})}$, $\gamma=\langle\hat{U}_{1}^{*},\mathbf{Q}_{0,0}\rangle_{2}$ and $y_{n}=r_{0}-\frac{n\pi}{4}$.
@@
We note that the core manifold $\mathcal{W}^{cu}_{-}(\mu)$ converges to a co-rotating two-dimensional manifold (3.21), with a co-rotating phase $\mathrm{e}^{\mathrm{i}(r_{0}-\frac{n\pi}{4})}$. We then expect the far-field manifold to possess a co-rotating phase $\mathrm{e}^{\mathrm{i}\phi(r)}$, where $\phi(r)$ is some perturbation from $(r-\frac{n\pi}{4})$; see Figure 3. We apply nonlinear normal form transformations to (3.20), as seen in [31, 22, 32], to remove the non-resonant terms from the right-hand side and the co-rotating phase $\mathrm{e}^{\mathrm{i}\phi(r)}$. For a more detailed proof of this result, see the analogous proof of Lemma 3.3 in Appendix A of [17].
@@
Having transformed our equations into the radial normal form (3.23) for an $n$-dimensional Hamiltonian–Hopf bifurcation, we also introduce the unconstrained variable $\kappa(r)$ to take the place of $\sqrt{\mu}$ so that the normal form (3.23) is extended to the following system,
@@
The methods presented in the rest of this section are analogous to several previous studies of localised radial patterns, and so we present only condensed proofs of each result. For more detailed exposition on this approach, see [18, 26, 22, 16, 17].
@@ 3.3 The Rescaling Chart
We now define rescaling coordinates in order to find exponentially decaying solutions for sufficiently large values of $r$; that is, we introduce
@@
These coordinates are the standard rescaling coordinates seen in [26]. Then, we can write (3.25) in the rescaling chart as,
@@
We present the following lemma regarding the existence of a nondegenerate solution to (3.31), which is required for the existence of ring and spot B solutions. We hence impose the restriction $40<n<$ 4, as required for such solutions in Therorem 2.8.
@@
We begin by noting that bounded localised solutions of (3.31) must also satisfy
$\begin{split}0={}&\int_{0}^{\infty}A_{R}(s)\left\{-\mathcal{D}_{\frac{n}{2}}\mathcal{D}_{\frac{n}{2}}A_{R}(s)+c_{0}A_{R}(s)+c_{3}|A_{R}(s)|^{2}A_{R}(s)\right\}\,s^{n}\,\mathrm{d}s\\ ={}&\int_{0}^{\infty}\left\{|\mathcal{D}_{\frac{n}{2}}A_{R}(s)|^{2}+c_{0}|A_{R}(s)|^{2}+c_{3}|A_{R}(s)|^{4}\right\}\,s^{n}\,\mathrm{d}s,\end{split}$
and so we see that either $c_{3}<0$ or $A_{R}\equiv 0$. Assuming that $c_{3}<0$, we note that finding forward-bounded localised solutions to (3.31) is equivalent to finding radially-symmetric ground state solutions to
$\Delta\hat{u}=\hat{u}-|\mathbf{x}|^{2-n}\hat{u}^{3},\qquad\hat{u}(\mathbf{x})\in\mathbb{R},\qquad\mathbf{x}\in\mathbb{R}^{3}.$ (3.33)
@@
which follows from the identities presented in (2.6). Hence the existence and non-degeneracy of $\hat{q}(s)$ follows from Proposition 2.7. Finally, the asymptotic behaviour of $q(s)$ in (3.32) can be found by noting that $q(s)$ is a solution of the rescaled equation
@@
We conclude this subsection by noting that upon inverting the rescaling transformation (3.27), exponentially decaying solutions to (3.25) evaluated at $r=r_{1}\mu^{-\frac{1}{2}}$ either take the form,
@@ 3.4 The Transition Chart
In the previous subsection we parametrised the set of exponentially decaying solutions to (3.25) for $r>r_{1}\mu^{-\frac{1}{2}}$. Now, in order to match these exponentially decaying solutions with the core manifold described in §3.1, we must track the trajectories associated with (3.40) and (3.41) backwards through the ‘transition’ region $r_{0}\leq r\leq r_{1}\mu^{-\frac{1}{2}}$. This leads to the following result.
@@
For each fixed choice of $0<r_{1},r_{0}^{-1}\ll 1$, there are constants $a_{0},\kappa_{0}>0$ such that solutions of (3.25), evaluated at $r=r_{0}$, $\kappa(r_{0})=\mu^{\frac{1}{2}}$, for initial values (3.40-3.41) are given by the following forms for all $\mu<\kappa_{0}$, $|a|<a_{0}$ and $Y\in\mathbb{R}$.
@@
We begin by solving (3.25) for $\kappa(r)$ and $\sigma(r)$ explicitly, giving
@@
For each initial value (3.40, 3.41), we utilise this approach in conjunction with appropriate choices of $R_{-}$, $R_{+}$ and coordinate transformations for $(A,B)$ in order to obtain our expressions for $(A(r_{0}),B(r_{0}))$.
@@
Inverting the transformations (3.46,3.48) brings us to the desired result. ∎
@@ 3.5 Matching Core and Far-field Manifolds
In the previous subsection we have obtained a parametrisation for exponentially-decaying trajectories of (3.18), which lie on the far-field manifold $\mathcal{W}^{s}_{+}(\mu)$, evaluated at the fixed point $r=r_{0}$. We now seek intersections of these trajectories with the parametrisation of the core manifold $\mathcal{W}^{cu}_{-}(\mu)$, evaluated at $r=r_{0}$, defined in (3.17).
We begin by applying (3.22) to the core parametrisation (3.21), such that the core manifold $\mathcal{W}^{cu}_{-}(\mu)$ can be expressed as
@@
where we recall $\nu_{n}=\left(\frac{3}{8}\right)^{\frac{n}{2}}\frac{\pi}{3\Gamma(\frac{n}{2})}$, $\gamma=\langle\hat{U}_{1}^{*},\mathbf{Q}(\hat{U}_{0},\hat{U}_{0})\rangle_{2}$ and introduce $Y_{n}=-\frac{n\pi}{4}+\mathcal{O}(|\mu|r_{0}+r_{0}^{-2})$.
@@
For each fixed choice of $0<r_{1},r_{0}^{-1}\ll 1$, there is a constant $\kappa_{0}>0$ such that intersections between the core manifold (3.52) and respective far-field trajectories (3.42-3.44) satisfy the following conditions for all $\mu<\kappa_{0}$.
@@
Having determined the leading order solution in each case, we now apply the implicit function theorem to solve (3.59, 3.61, 3.65, 3.68) uniquely for all $0<\mu,r_{1},r_{0}^{-1}\ll 1$. Inverting the respective coordinate transformations (3.58, 3.60, 3.64, 3.67), we arrive at our result. ∎
@@ 4 Discussion
In this paper, we have proven that localised $(n+1)$-dimensional radial patterns emerge from non-degenerate Turing bifurcations in general two-component reaction-diffusion systems. Extending the approach of [22, 26], we determined that spot A patterns emerge for all values of the dimension parameter $n>0$, where the core profile depends continuously on $n$ and possesses a fixed asymptotic scaling $\mu^{\frac{1}{2}}$. In contrast, we found that ring and spot B patterns only emerge for $40<n<$ 4, where their core profiles depend continuously on $n$ and they possess an asymptotic scaling $\mu^{\frac{4-n}{4}}$, $\mu^{\frac{4-n}{8}}$, respectively. The core profile for each pattern depends on $(n+1)$-dimensional Bessel functions $J^{(n)}_{\ell}(r)$, which can be written in terms of standard Bessel functions; this highlights the connection between core profiles of two-dimensional and three-dimensional radial patterns that was previously unclear. Our results agree with previous findings in [22, 26], while providing a proof for the previously unproven existence of three-dimensional rings and presenting simplified proofs for spot B patterns and solutions of the non-autonomous Ginzburg–Landau equation (3.30).
@@
While we have considered the emergence of localised radial patterns from an $(n+1)$-dimensional Turing instability, very little is understood of their behaviour away from this bifurcation point. As seen in Figure 1(a), spot A patterns may undergo snaking behaviour as they move away from their bifurcation point. This is behaviour is known as homoclinic snaking (first introduced in [40]) when $n=0$, which has been the subject of over two decades of research, (see [9]) but is still not well-understood for $n>0$. While there has been some numerical studies of the bifurcation structure of spots in two and three dimensions [25, 13], and analytical results regarding ring patterns in $n$-spatial dimensions [6], this topic still requires further investigation.
A natural next step in this analysis would be to consider the stability of localised radial patterns for various spatial dimensions. In particular, the transition of localised one-dimensional patterns to axisymmetric patterns (and from axisymmetric to spherically-symmetric) would be an interesting phenomenon to explore. Our current approach could also be extended to higher-dimensional symmetries, such as the dihedral patterns considered in [16, 17]. This would require some kind of formulation of $(n+1)$-dimensional spherical harmonics that vary continuously with respect to $n$. Being able to describe the transition of more general patterns into their higher- or lower-dimensional counterparts would an extremely powerful tool for understanding pattern formation in nature.
  • Acknowledgments. The author gratefully acknowledges support from the Alexander von Humboldt foundation, and would like to thank David Lloyd for his comments on an earlier draft of this manuscript.

### v3

@@ 3.2 The Far-Field Manifold
We seek to find all small-amplitude exponentially-decaying solutions to (2.1), thereby constructing the far-field manifold $\mathcal{W}_{+}^{s}(\mu)$. We introduce the variable $\sigma(r)\geq 0$ to replace any $\frac{1}{r}$ terms in (3.1), so that we obtain the autonomous extended system
@@
with the property that $\sigma(r)=r^{-1}$ is an invariant manifold of (3.17). This invariant manifold clearly recovers the non-autonomous system (3.1), and so any solutions to (3.17) where $|\mathbf{U}|\to 0$ exponentially fast as $r\to\infty$ and $\sigma(r_{0})=r_{0}^{-1}$ is an exponentially decaying solution of (3.1).
We first transform (3.17) into a radial normal form so that we can identify exponentially decaying solutions. We introduce complex amplitudes $\widetilde{A},\widetilde{B}$, where
@@
The core manifold (3.16) can be expressed as
@@
where $\nu_{n}=\left(\frac{3}{8}\right)^{\frac{n}{2}}\frac{\pi}{3\Gamma(\frac{n}{2})}$, $\gamma=\langle\hat{U}_{1}^{*},\mathbf{Q}_{0,0}\rangle_{2}$ and $y_{n}=r_{0}-\frac{n\pi}{4}$.
@@
We note that the core manifold $\mathcal{W}^{cu}_{-}(\mu)$ converges to a co-rotating two-dimensional manifold (3.20), with a co-rotating phase $\mathrm{e}^{\mathrm{i}(r_{0}-\frac{n\pi}{4})}$. We then expect the far-field manifold to possess a co-rotating phase $\mathrm{e}^{\mathrm{i}\phi(r)}$, where $\phi(r)$ is some perturbation from $(r-\frac{n\pi}{4})$; see Figure 3. We apply standard normal form transformations, as seen in [35, 26, 36], to (3.19) so that all low-order non-resonant terms and the co-rotating phase $\mathrm{e}^{\mathrm{i}\phi(r)}$ are removed from the right-hand side. For a more detailed proof of this result, see the analogous proof of Lemma 3.3 in Appendix A of [20].
@@
The resultant system (3.22) is the radial normal form for an $n$-dimensional Hamiltonian–Hopf bifurcation, which we extend further by introducing the unconstrained variable $\kappa(r)$ to take the place of $\sqrt{\mu}$. Then, (3.22) becomes
@@
The methods presented in the rest of this section are analogous to several previous studies of localised radial patterns, in particular the recent papers [19, 20], and so we present only condensed proofs of each result. For more detailed exposition on this approach, see [21, 30, 26, 19, 20].
@@ 3.3 The Rescaling Chart
To capture the ‘slow’ behaviour of (3.24) for large values of $r$, we introduce the standard rescaling coordinates of [30]
@@
Then, we can write (3.24) in the rescaling chart as,
@@
We present the following lemma regarding the existence of a nondegenerate solution to (3.30), which is required for the existence of ring and spot B solutions. We hence impose the restriction $40<n<$ 4, so that we can apply Proposition 2.7 and Hypothesis 2.8, as required for such solutions in Theorem 2.9.
@@
We note that finding forward-bounded localised solutions to (3.30) is equivalent to finding radially-symmetric ground state solutions to
$\Delta\hat{u}=\hat{u}+\mathrm{sgn}\,(c_{3})|\mathbf{x}|^{2-n}\hat{u}^{3},\qquad\hat{u}(\mathbf{x})\in\mathbb{R},\qquad\mathbf{x}\in\mathbb{R}^{3}.$ (3.32)
@@
which follows from the identities presented in (2.6). A standard variational argument shows that bounded localised solutions must also satisfy
$\begin{split}0={}&\int_{\mathbb{R}^{3}}\hat{u}\left(-\Delta\hat{u}+\hat{u}+\mathrm{sgn}\,(c_{3})|\mathbf{x}|^{2-n}\hat{u}^{3}\right)\,\mathrm{d}\mathbf{x}=\int_{\mathbb{R}^{3}}\left\{|\nabla\hat{u}|^{2}+|\hat{u}|^{2}+\mathrm{sgn}\,(c_{3})|\mathbf{x}|^{2-n}|\hat{u}|^{4}\right\}\,\mathrm{d}\mathbf{x},\\ \end{split}$
and so we see that either $c_{3}<0$ or $\hat{u}\equiv 0$. Hence the existence and non-degeneracy of $\hat{q}(s)$ follows from Proposition 2.7 for $30<n<$ 3 and Hypothesis 2.8 for $43\leq n<$ 4. Finally, the asymptotic behaviour of $q(s)$ in (3.31) can be found by noting that $q(s)$ is a solution of the rescaled equation
@@
After inverting the rescaling transformation (3.26), exponentially decaying solutions to (3.24) evaluated at $r=r_{1}\mu^{-\frac{1}{2}}$ then either take the form
@@ 3.4 The Transition Chart
So far, we have identified exponentially decaying trajectories in (3.24) for $r>r_{1}\mu^{-\frac{1}{2}}$. To connect these trajectories with the core manifold constructed in §3.1, we must track solutions with boundary data (3.39) and (3.40) backwards through the ‘transition’ region $r_{0}\leq r\leq r_{1}\mu^{-\frac{1}{2}}$ to the boundary of the core manifold at $r=r_{0}$.
@@
For each fixed choice of $0<r_{1},r_{0}^{-1}\ll 1$, there are constants $a_{0},\kappa_{0}>0$ such that solutions of (3.24), evaluated at $r=r_{0}$, $\kappa(r_{0})=\mu^{\frac{1}{2}}$, for initial values (3.39-3.40) are given by the following forms for all $\mu<\kappa_{0}$, $|a|<a_{0}$ and $Y\in\mathbb{R}$.
@@
We note that the equations for $\kappa$ and $\sigma$ in (3.24) decouple from $A$ and $B$, and so we begin by solving (3.24) for $\kappa$ and $\sigma$ with initial data (3.39-3.40), giving
@@
For each initial value (3.39, 3.40), we utilise this approach in conjunction with appropriate choices of $R_{-}$, $R_{+}$ and coordinate transformations for $(A,B)$ in order to obtain our expressions for $(A(r_{0}),B(r_{0}))$.
@@
Inverting the transformations (3.45, 3.47) brings us to the desired result. ∎
@@ 3.5 Matching Core and Far-field Manifolds
In the previous subsection we have obtained a parametrisation for exponentially-decaying trajectories of (3.17), which lie on the far-field manifold $\mathcal{W}^{s}_{+}(\mu)$, evaluated at the fixed point $r=r_{0}$. We now seek intersections of these trajectories with the parametrisation of the core manifold $\mathcal{W}^{cu}_{-}(\mu)$, evaluated at $r=r_{0}$, defined in (3.16). We recall from Remark 3.2 that the core manifold $\mathcal{W}^{cu}_{-}(\mu)$, following the normal form transformations (3.21) can be written as
@@
where $\nu_{n}=\left(\frac{3}{8}\right)^{\frac{n}{2}}\frac{\pi}{3\Gamma(\frac{n}{2})}$, $\gamma=\langle\hat{U}_{1}^{*},\mathbf{Q}(\hat{U}_{0},\hat{U}_{0})\rangle_{2}$ and $Y_{n}=-\frac{n\pi}{4}+\mathcal{O}(|\mu|r_{0}+r_{0}^{-2})$.
@@
For each fixed choice of $0<r_{1},r_{0}^{-1}\ll 1$, there is a constant $\kappa_{0}>0$ such that intersections between the core manifold (3.51) and respective far-field trajectories (3.41-3.43) satisfy the following conditions for all $\mu<\kappa_{0}$.
@@
Having determined the leading order solution in each case, we now apply the implicit function theorem to solve (3.58, 3.60, 3.64, 3.67) uniquely for all $0<\mu,r_{1},r_{0}^{-1}\ll 1$. Inverting the respective coordinate transformations (3.57, 3.59, 3.63, 3.66), we arrive at our result. ∎
@@ 4 Discussion
In this paper, we have proven that localised $(n+1)$-dimensional radial patterns emerge from non-degenerate Turing bifurcations in general two-component reaction-diffusion systems. Extending the approach of [26, 30, 19, 20], we determined that spot A patterns emerge for all values of the dimension parameter $n>0$, where the core profile depends continuously on $n$ and possesses a fixed asymptotic scaling $\mu^{\frac{1}{2}}$. In contrast, we found that ring and spot B patterns only emerge for $40<n<$ 4, where their core profiles depend continuously on $n$ and they possess an asymptotic scaling $\mu^{\frac{4-n}{4}}$, $\mu^{\frac{4-n}{8}}$, respectively. The core profile for each pattern depends on $(n+1)$-dimensional Bessel functions $J^{(n)}_{\ell}(r)$, which can be written in terms of standard Bessel functions; this highlights the connection between core profiles of two-dimensional and three-dimensional radial patterns that was previously unclear. We observe that spots are larger and sharper as $n$ increases, and that their oscillations are more strongly damped. This suggests that it may be harder to identify the associated wave number of localised three dimensional spots in physical and biological examples, preventing their characterisation as Turing patterns. However, our results show that such patterns are strongly associated with Turing bifurcations as well as their lower-dimensional counterparts. Our results agree with previous findings in [26, 30], while providing a proof for the previously unproven existence of three-dimensional rings and presenting simplified proofs for spot B patterns and solutions of the non-autonomous Ginzburg–Landau equation (3.29) with $30<n<$ 3.
@@
While we have considered the emergence of localised radial patterns from an $(n+1)$-dimensional Turing instability, very little is understood of their behaviour away from this bifurcation point. As solutions become larger, the nonlinearity of (1.2) becomes more important and localised radial solutions may no longer reflect the profiles predicted by our analysis. As seen in Figure 1(a), spot A patterns may undergo snaking behaviour as they move away from their bifurcation point. This is behaviour is known as homoclinic snaking (first introduced in [45]) when $n=0$, which has been the subject of over two decades of research, (see [8]) but is still not well-understood for $n>0$. While there has been some numerical studies of the bifurcation structure of spots in two and three dimensions [29, 14, 38], and analytical results regarding ring patterns in $n$-spatial dimensions [5], this topic still requires further investigation.
A natural next step in this analysis would be to consider the stability of localised radial patterns for various spatial dimensions, for which there has been a recent numerical study [38]. In particular, the transition of localised one-dimensional patterns to axisymmetric patterns (and from axisymmetric to spherically-symmetric) would be an interesting phenomenon to explore. A rigorous understanding of the stability of localised radial patterns is not currently well established, even in the simple cases of $n=1,2$, and so this would be a fruitful area of future study. Our current approach could also be extended to higher-dimensional symmetries, such as the dihedral patterns considered in [19, 20]. Much of the approach presented § 3 is equivalent to the results presented in [19, 20], and our simplified proof for Spot B patterns without blow-up coordinates presents a new strategy for obtaining dihedral spot B-type patterns in a future work. However, studying the dimensional effects of fully localised non-radial patterns would require some kind of formulation of $(n+1)$-dimensional spherical harmonics that vary continuously with respect to $n$ and would be a highly nontrivial extension of the work presented in [19, 20].
  • Acknowledgments. The author gratefully acknowledges support from the Alexander von Humboldt foundation, and would like to thank David Lloyd for his comments on an earlier draft of this manuscript. Additionally, the author would like to thank Boris Buffoni and François Genoud for their valuable insight on ground states to elliptic PDEs.